adios. The direct During the mosth of 13. Zuany Electronies Coeporation uses a standard cost system for the peoduction of s ts water ski labor standard for each radio is 0.9 hours. The standand direct labor cost per hour is 57.0 Asgust, Zannys water ski radio What is Zarnys labor rate variance for August? prodection used 6,600 direct labor-hours at a total direct labor cost of $48,708. a$972 Favorable b. $1,188 Unfavorable c. $2,160 Faverable d. $2,808 Unfavorable 14. Which of the following would resulk in a labor rate variance: Poor quality of raw materials b. Replacing a skilled work force will unskilled labor c. Excessive overhead spending d. Employee poor asitude toward management 15. Which of the fhollowing Cooting System? companies is more likely to use a Process Costing System versus and Job-Oeder a. Apple Computers b. Van Husen Shirt Company e Rolex Watches d. Coca-Cola Case Coporationproduces and sellsasingleprodat. The selling pree is S 170, per unit and the product variable cost is 568 per unit. units per month. Management is considering using a new component that would increase the unit variable cost by T Fixed expenses are $521,000 per month. The company is currently selling 7,000 $6. Since the new coemponent would increase the features of the companys product, the marketing manager predicts that monthly sales would incrcase by 500 units. Irall works out as expected, what should be the overall effect on the companys monthly net operating income of this change? 17. Windstorm Corporation produces and sells a single product at a price of S130 per unit with a variable cost per unit of $41.60. How many units do they need to sell next month and what is their roquired total sales dollars to cover their fixed costs of $109,616 18. The Krewe of Students is having a Bingo Toumament to raise moncy for a family in-need. Tbeir pay-out to the winners for the evening is $1,500. They are renting the convention center for $500 and paying the bimgo announcer S100. Their goal for is to donate $5,000 to the family. They will be selling the bingo packet at $35 each and each packet costs the Krewe $2.50 a. How many bingo packets do they need to sell to reach their goal? b. A krewe member said that they would sell more packets if they dropped the sales price of each packet to $25 and gave a glass of wine for each packet sold. Each glass of wine costs $2.00. How many MORE packets would they have to sell to reach their goal?
Answer #1

13)option:B

Labour rate variance:

It shall be the change in the amount of labour rate per standard hours.

Hence cost of 6600 labour hours @7.2per hour shall be 47520.

Hence the cost to company is 48708.

Hence labour rate variance shall be47520-48708=1188(unfavourable).

14)option:B

labour rate variance shall be the no.of replacements &removals.

It shall be said as the change in labour .

Hence it shall be the replacement of unskilled labour in place of skilled labour.

15)Option:B

As the shirt company has a specific job as per the measurements of individuals and each one shall have a each job order and undergoes a process of different departments.

16)

Sales shall increase by 500.hence sales value shall be

7500*170=1275000

Variable costs shall be 68*7500=(510000)

Fixed costs shall be(521000)

Hence net income =1275000-510000-521000=244000

Hence effect of change is 193000(given as per information).

Hence effect is 244000-193000=51000

Note: As per HOMEWORKLIB POLICY,we need to awnser the first four internal questions if the internal questions are more than 4.hence cooperate.